### v1

# The ML Supply Chain in the Era of Software 2.0: Lessons Learned from Hugging Face

Abstract.
The last decade has seen widespread adoption of Machine Learning (ML) components in software systems. This has occurred in nearly every domain, from natural language processing to computer vision. These ML components range from relatively simple neural networks to complex and resource-intensive large language models. However, despite this widespread adoption, little is known about the supply chain relationships that produce these models, which can have implications for compliance and security. In this work, we conduct an extensive analysis of 760,460 models and 175,000 datasets mined from the popular model-sharing site Hugging Face. First, we evaluate the current state of documentation in the Hugging Face supply chain, report real-world examples of shortcomings, and offer actionable suggestions for improvement. Next, we analyze the underlying structure of the extant supply chain. Finally, we explore the current licensing landscape against what was reported in prior work and discuss the unique challenges posed in this domain. Our results motivate multiple research avenues, including the need for better license management for ML models/datasets, better support for model documentation, and automated inconsistency checking and validation. We make our research infrastructure and dataset available to facilitate future research.
1. Introduction
The use of machine learning (ML) models in software applications has increased dramatically over the last decade, including in recommendation systems, computer vision, chatbots, image generation, automated software engineering, and more. As creating and training new models has become increasingly more expensive (Patterson et al., 2021), developers have turned to fine-tuning pre-existing models. This approach may save time and effort, but it also introduces the complexity of a new ML supply chain, complete with many novel challenges.
While supply chains for conventional software typically consist of software components, libraries, configuration files, and processes (Tan et al., 2022; Gao et al., 2024), for ML-intensive systems, the supply chain is more complex, comprising all elements that make an ML algorithm work. Creating a modern ML system involves the integration of ML models and traditional software components, as well as different types of reuse or integration, taking into account dependencies between models themselves or between models and their training datasets. In addition, the training of ML models can rely on multiple datasets or dataset aggregates, and datasets, in turn, each have their own supply chain (Lee et al., 2023).
Adaptation or reuse of existing models can also introduce degrees of complexity. Models can be fine-tuned (Church et al., 2021) or quantized from existing models (Gholami et al., 2022). (Fine-tuning allows a developer to use an existing model to create a model that specializes in certain tasks or to change the model’s behavior, such as removing safety or ethical protections. Quantization compresses an existing model such that it can be run on smaller hardware with minimal performance loss.) Different ML models can be combined to form a single architecture that can, in turn, be reused or adapted (Ardabili et al., 2019). Finally, the outputs from one model can be used to train another model, such as through synthetic data (Marwala et al., 2023) or the use of a teacher-student approach (Hinton, 2015).
The ML supply chain, particularly those involving generative models (Lee et al., 2023), is not necessarily linear, progressing cleanly from one dependency or stage to the next. Instead, as outlined by Lee et al. (Lee et al., 2023), there can be branches and even cycles in the supply chain. For example, datasets can be used not only in the initial training of models but also for the fine-tuning of pre-existing base models, sometimes in multiple instances and by different developers. Additionally, some stages of the supply chain can backfeed into others, making relationships increasingly complex, if not recursive. For example, the outputs of a generative model can be added to pre-existing datasets and used for training future versions. The sheer volume of content from generative models makes this situation increasingly likely, if not inevitable.
Understanding these relationships is critical not only for license compliance, which requires a full understanding of the components used in a project and their associated licenses, but also for detecting, mitigating, and managing security threats involved with model reuse, such as weight poisoning attacks (Kurita et al., 2020), data poisoning (Goldblum et al., 2020), and even malware hidden in model weights and assembled at runtime (Wang et al., 2021).
While previous work has considered the documentation (Pepe et al., 2024), evolution (Jiang et al., 2022), and environmental impact (Castaño et al., 2024a) of ML models, the complexity and challenges of the ML supply chain as a whole have not yet been fully explored. The supply chain relationships between models and dependent GitHub repositories have been explored by Pepe et al. (Pepe et al., 2024) and Jiang et al. (Jiang et al., 2024b), but the relationships between ML models remain understudied.
To contribute to bridging this gap, we aim to investigate the emerging ML model supply chain on Hugging Face (Inc., 2024) (hereafter referred to as the ‘ML supply chain’), specifically the documentation practices of model owners, the structure and complexity of the supply chain itself, and the existence and prevalence of potential compliance issues. We note that proper documentation and documentation practices are foundational to mapping, and thus managing the ML supply chain. Launched in 2016, Hugging Face is, as of January 2025, the largest repository of ML models and datasets (Jiang et al., 2023c, 2022). As with GitHub and traditional software, Hugging Face provides a central hub for model developers and data scientists to explore, share, and experiment with ML models. While other model forges—such as TensorFlow Hub (Google Inc., 2024), PyTorch Hub (The Linux Foundation, 2024), and Model Zoo (Koh, 2024)—exist, the Hugging Face platform has the widest reach, hosting more than 750K models and 175K datasets across many different use cases. Hugging Face therefore represents a robust opportunity to understand ML supply chains and analyze how developers interact with them.
We mine models and datasets hosted on Hugging Face, using the available information to construct a supply chain graph. We highlight different types of documentation inconsistencies that complicate the creation of this supply chain graph, discuss the underlying structure and features of the current ML supply chain on Hugging Face, and analyze the present licensing landscape. Ultimately, we demonstrate that there is still much work to be done to facilitate better dependency/supply chain management on Hugging Face and in the ML ecosystem.
The main contributions of this paper are the following:
  • A critical evaluation of the current state of documentation on Hugging Face.
  • An analysis of the interactions between models and datasets in the supply chain.
  • An analysis of the licensing ecosystem on Hugging Face and potential compliance problems.
  • Actionable suggestions for how model hubs can better facilitate supply chain management activities.
  • A cleaned, reusable, and extensible dataset of models, datasets, and their relationships, and a tool suite capable of generating newer versions of said dataset (Stalnaker et al., 2025).
2. Background and Related Work
The ML supply chain has emerged as a complex structure with many steps and components, each of which can affect the final output of a given model. Lee et al. identify eight stages in the generative AI supply chain, but most apply more broadly: 1) the creation of expressive works that will eventually be used to train a model, 2) the conversion of these expressive works into digital data, 3) the compilation of these data points into training datasets, 4) the creation of an ML model by selecting an architecture, training datasets, and a training algorithm, 5) the fine-tuning of existing “base” models, 6) the deployment of the trained/fine-tuned model as a service, 7) the use of the model to generate output, and finally 8) applying additional alignment to further improve the model or meet user needs (Lee et al., 2023). Modern ML models are also becoming increasingly complex with respect to their architectures. For example, Mixture of Expert (MoE) models and ensemble models comprise multiple specialized models trained by dividing the problem space into homogeneous regions (Masoudnia and Ebrahimpour, 2014; team, 2023). Recent work has also proposed using smaller ML models to replace the hidden states within a Recurrent Neural Network (RNN) architecture to facilitate better recall and compression of context (Sun et al., 2024). Both of these examples demonstrate that understanding the relationships among models and among model architectures will become increasingly difficult.
As a result, sharing information about models has become critically important. ML components do not, however, always disclose data sources, making it difficult to comply with or even keep track of licensing obligations associated with the data (Hassan et al., 2024). This creates conditions that could lead to legal disputes over the use of copyrighted material in training ML components (Vincent, 2022). Model cards (Mitchell et al., 2019) have become the standard method for sharing information about ML models hosted on Hugging Face (Face, ndb), including intended uses, limitations, and datasets used in training (Face, nda). Although tools are in development to streamline the process (Face, ndc), model card creation on Hugging Face is primarily a manual process in which a user enters documentation about a model into a predefined markdown template (Face, 2024b). This template is robust, including spaces for model description, uses, bias, limitations, testing, etc., as well as optional fields for citation, technical specifications, and environmental impact. The manual nature of documentation, however, introduces a significant potential for human error, ambiguity, and incompleteness (Jiang et al., 2024b, 2023d). Indeed, the quality and adoption of model cards remains low (Bhat et al., 2023), despite user studies and other efforts by Hugging Face (Face, [n. d.]) to ameliorate the situation, and despite researchers proposing tools such as DocML (Bhat et al., 2023) to create and evolve model cards.
Hugging Face has previously been mined to study the maintenance and evolution of ML models (Castaño et al., 2024b), security risks in the ML supply chain (Jiang et al., 2022), the carbon footprint of ML models (Castaño et al., 2023), and the documentation of ML models in terms of datasets, bias, and licenses (Pepe et al., 2024). Through an analysis of discussion forums and a mining study of Hugging Face, Taraghi et al. (Taraghi et al., 2024) identified challenges and benefits associated with model reuse. Also, previous work identified challenges and deficiencies in the organization of the ML supply chain, including poorly chosen or erroneous model names (Jiang et al., 2023a) and frequently missing or incorrect licensing information (Longpre et al., 2023) on Hugging Face and other platforms.
While prior work has previously noted the incompleteness of model/dataset documentation on Hugging Face (Jiang et al., 2023c; Pepe et al., 2024; Yang et al., 2024), our work goes beyond this observation and provides a detailed analysis of specific examples in which documentation is incomplete and erroneous. These real-world observations can help model owners avoid common documentation pitfalls, empower model hubs to implement automated checks for common errors, and make researchers aware of potential noise in data mined from Hugging Face.
Oreamuno et al. (Oreamuno et al., 2024) also observed that the majority of models and datasets on Hugging Face lacked documentation. When present, this documentation was in the form of model/data cards, was often incomplete, and followed standards inconsistently. Whereas the work of Oreamuno et al. focused on exploring the presence and completeness of documentation, we report various inconsistencies and anomalies that are introduced by Hugging Face users during the process of documenting their models and datasets. The anomalies we encountered can exist even in cases where the model/dataset documentation is complete.
Yang et al. (Yang et al., 2024) conducted an in-depth analysis of the state of dataset documentation on Hugging Face, focusing primarily on the content included in the documentation (e.g., description, structure, creation). Our study expands on this work by also considering the documentation of ML models in addition to datasets and by noting errors or inconsistencies introduced by model/dataset owners.
Jiang et al. (Jiang et al., 2024a) studied naming practices and naming anomalies for models on Hugging Face. They found that model names carry more semantic information than do the names of traditional software packages, so anomalies, such as misrepresenting the model architecture, can lead to greater confusion in the ML context. Our work does not consider the accuracy of pre-trained model (PTM) names when compared against the metadata they provide; rather we note difficulties in constructing the ML supply chain given naming inconsistencies and ambiguities in metadata references.
Pepe et al. (Pepe et al., 2024) conducted the study most closely related to ours. However, instead of focusing on the structure of the ML supply chain, they explore the extent to which models document datasets, potential biases, and licenses. Both Pepe et al. (Pepe et al., 2024) and Jiang et al. (Jiang et al., 2024b) analyze a supply chain, but they focus on the relationship between models and downstream GitHub repositories that make use of those models. We instead focus on the relationships among PTMs, dependent models, and datasets.
Both Pepe et al. (Pepe et al., 2024) and Jiang et al. (Jiang et al., 2024b) also explore model licensing on Hugging Face. In the case of Jiang et al. (Jiang et al., 2024b), this is done as a short demonstration illustrating the utility of the PEATMOSS dataset. We extend these prior works by collecting and analyzing more recent data, which reflects the current state of model/dataset licensing. Both works also consider the interplay of model licenses with the licenses of dependent GitHub repositories. Neither work, however, evaluates license incompatibilities between models or between models and datasets. We further expand on prior work by considering various license classes, multi-licensing, and the failure to comply with license-imposed naming requirements. Where previous work examines the relationships between projects and models, we examine the relationships between multiple models. Thus, our work is useful not only to software developers, as is the case for prior work, but also to data scientists, those who want to reuse models that have complex dependencies, and those who want to develop new models. By providing an ecosystem-level examination of the ML supply chain, we provide these groups with the information necessary to understand how models relate to each other and evolve, licensing concerns regarding ML components on Hugging Face, and potential challenges that might be encountered regarding model documentation.
In conclusion, while our study has been informed by prior work (Castaño et al., 2024a; Yang et al., 2024; Jiang et al., 2024a, 2023d; Oreamuno et al., 2024; Jiang et al., 2023c; Pepe et al., 2024; Taraghi et al., 2024; Jiang et al., 2023b), our investigation focuses on the structure and characteristics of the ML supply chain itself and the challenges encountered in managing it. More specifically, while previous work examined deficiencies in documentation and licensing at the level of individual models, we investigate the ML supply chain at a higher level by examining how it is structured and how models relate to each other.
3. Study Methodology
The goal of this study was to evaluate the current state of documentation, structure, and licensing of the ML supply chain. The study context consisted of 760,460 models and 175,000 datasets mined from Hugging Face. We aim to address the following research questions (RQs):
  1. RQ0:
    What documentation deficiencies exist in Hugging Face model and data cards that potentially complicate mapping the ML supply chain? To understand the ML supply chain, we needed to collect and analyze Hugging Face metadata. Using this data, we identified several issues and challenges related to documentation that introduced hurdles to obtaining a comprehensive understanding of the ecosystem. We examine these challenges in detail because with poorly and inconsistently described models and datasets, developers may struggle to leverage the ML supply chain effectively. This preliminary research question discusses encountered examples of noisiness issues occurring in Hugging Face model and dataset documentation.
  2. RQ1:
    What is the structure of the ML supply chain? The ML supply chain, like any other supply chain, is characterized by dependencies across components. This research question looks into the complexities of such a dependency graph.
  3. RQ2:
    What is the current licensing landscape for models and datasets and what potential compliance challenges does it pose? ML models are software, so software licenses govern their use and redistribution. As such, these models present new challenges in the area of license compliance. To understand these challenges, we examine the state of licensing in the ML supply chain by 1) analyzing common licenses and license categories, 2) exploring licensing decisions made in the supply chain, and 3) investigating potential compliance challenges, such as incompatibilities between parent/child model licenses, ambiguities introduced by undeclared, unknown, or custom licenses, and uncertainty surrounding how license terms apply in context.
Figure 1. Study Methodology
In the following subsections, we describe the study methodology, which is depicted in Fig. 1. We start by mining model and dataset lists. Next, for each model/dataset, we extract information necessary to construct the ML supply chain graph, recording and analyzing documentation deficiencies encountered during the process (RQ0). Then, we systematically analyze and evaluate the structure of the resulting supply chain graph (RQ1). This same graph, along with licensing information, is then used to address RQ2.
3.1. Data collection
To answer the research questions, we used the Hugging Face API to extract a list of all models publicly hosted on the platform. An initial list of 500,000 models was pulled on July 9, 2024, and a full list of all available models was obtained on July 11, 2024. Using the same APIs, we mined data (model cards and metadata) for each model on this list. For models that were unavailable through the API (as described in Section 4.1), we employed a web-scraping technique relying on a combination of Python’s request module and the BeautifulSoup HTML parser. Mining took place between July 9 and July 12, 2024, yielding 760,460 mined models. We also used the Hugging Face API to gather the list of all datasets publicly hosted on the platform and available through the API as of July 9, 2024. We make all the raw data acquired from mining available in our replication package (Stalnaker et al., 2025).
While a handful of research datasets for Hugging Face models already exist (Jiang et al., 2024b, 2023d), we opted to create our own dataset for this work for three reasons. First, we aimed to analyze the most current data, which includes new models and model families (e.g., new models recently added to the Llama family of models) and changes in licensing frequencies (see Section 6). Second, we aimed to collect and analyze the most comprehensive and complete dataset; as shown in Table 1, our dataset is significantly larger than existing datasets. Third, unlike Jiang et al. (Jiang et al., 2024b) and Pepe et al. (Pepe et al., 2024), we were not focused on model-to-GitHub repository dependencies; accordingly, we aimed to design a dataset tailored for mapping model-to-model dependencies.
Table 1. Comparison of Hugging Face (HF) datasets used in prior work
Work Models Datasets Study Goal(s)
(Yang et al., 2024) - 24,065 Explore the state of dataset documentation on HF
(Jiang et al., 2024a) 500 - Investigate PTM* naming practices/anomolies on HF
(Jiang et al., 2023d) 15,913 - Create a dataset for future research
(Oreamuno et al., 2024) 55,280 6,758 Understand the state of model documentation practices
(Jiang et al., 2023c) 63,182 - Understand model reuse, developer workflow, and challenges
(Pepe et al., 2024) 159,132 - Study extent to which model maintainers document PTMs
(Taraghi et al., 2024) 239,422 - Understand challenges faced during PTM reuse
(Jiang et al., 2023b) / (Jiang et al., 2024b) 281,638 - Create a dataset examining relationship between PTMs and OSS
(Castaño et al., 2024a) / (Castaño et al., 2024b) 380,000 - Provide guide for researchers/explore how models are maintained
This work 760,460 175,000 Explore structure of ML supply chain and licensing challenges
*Pre-Trained Model
3.2. Data normalization and cleaning
Once the data for models and datasets was downloaded, it needed to be cleaned, standardized, and stored in a more usable form. This process involved standardizing the names of declared base models and datasets and mapping to unique identifiers where possible. For example, xlm-roberta-base is a shorthand for FacebookAI/xlm-roberta-base, which has the unique internal ID 621ffdc036468d709f174364. We relied on Hugging Face’s internal mechanisms to resolve these differences where possible. For example, attempting to load the model page for xlm-roberta-base on Hugging Face redirects to the page for FacebookAI/xlm-roberta-base. Additionally, we resolved inconsistencies in field inputs. For example, [], “”, and None were all being used to denote no declared licenses. We standardized these variations to the empty list to make later computation simpler. We further elaborate on documentation challenges that required special attention and resolution in our discussion of RQ0.
3.3. Extracting licensing information
We extracted licensing information from the tags present in model metadata. While licenses can also be declared in the CardData attribute of the model metadata, we found through an analysis of our dataset that the tags were more complete in every instance than the CardData attribute. In nearly all instances, licensing information was in both locations (99.9%). However, for 0.1% of models, the licensing information was exclusively in the tags, and there were no instances that relied only on the CardData attribute. We also looked for discrepancies in those cases where both locations were utilized. We identified only 134 such discrepancies, and in all cases, an additional license was declared in the tags while it was not declared in the CardData. There were no irreconcilable differences (i.e., MIT vs GPL). From this, we concluded that the model tags are the most complete machine-readable source for licensing information.
Although licensing information can also be found in other, harder to programmatically extract sections of a model card, we sought to further confirm that reliance on tags to identify the declared license used by the model was justified. Accordingly, one author manually reviewed the model cards for the top 100 models (sorted by number of downloads) marked as having no license in the tags. We anticipated that these 100 models were more likely to have complete and reliable documentation because of their popularity as compared with other models in our dataset, consistent with the findings of Jones et al. (Jones et al., 2024). From this sample of 100 models, we identified only 16 models that had declared licensing information elsewhere. Of those, one had licensing information added to its tags after our initial mining, bringing the total to 15 models. For nine of these models, the license could only be found transitively by following a link to the model’s associated GitHub page. Table 2 provides a more detailed breakdown of where the licensing information was located for each example.
Table 2. License Information Location for Sample with No Tags after Manual Review
Location Count Percent
Link to GitHub 9 56.3%
License referenced in README 2 12.5%
License text embedded in README 2 12.5%
LICENSE File 2 12.5%
License in tags (added after mining) 1 6.3%
Total 16 100.0%
Detecting each model’s declared license would involve more complicated mining techniques, such as those that could traverse external links, and would require NLP-based approaches that extract licensing information from the text. These NLP approaches would have to not only recognize license names, which may not be in a standardized form but also pattern-match embedded license text. An LLM tool, like the one employed by Jiang et al. (Jiang et al., 2024b), could potentially be used to extract licensing information, but at a high cost (between \$0.01 and \$0.03 per model— or between \$7,000 and \$22,800 for our dataset). To ensure that our analysis captured the most accurate information possible and to better understand the formal documentation structure of Hugging Face, rather than employ such an approach, we focus on explicitly declared licenses. Given the likelihood that licenses are declared in tags, the infeasibility of manual annotation for all examples, and the inherent difficulties of automatically obtaining license information from anywhere in a model card, our reliance on tags for license detection appears to be an acceptable approach. Additionally, existing supply chain analysis tools rely heavily on machine-readable metadata, so we further justify that this is the appropriate locus of analysis given that the ultimate audience for such tools is the end user.
3.4. Data analysis
To address RQ0, we quantitatively and qualitatively analyzed and discussed the different types of documentation issues that we observed in our attempt to understand the ML supply chain on Hugging Face.
Once the dataset had been cleaned, to address RQ1, we used the networkx Python library (networkx, 2024) to create a directed graph representing the observable ML supply chain. Each node in the graph represents a distinct model, and an edge represents a dependency relationship. For example, if model $A$ listed model $B$ as a base model, there would be an edge from $B$ to $A$. We would describe model $B$ as the parent of model $A$ and model $A$ as the child of model $B$. Due to the inherent issues and incompleteness of model documentation, this graph, by its nature, only represents a subset of the total ML supply chain on Hugging Face.
For the sake of analysis, we flatten the individual model supply chains by extracting all longest paths (e.g., lineage chains) from the graph that go from source nodes to sinks. Each of these paths goes from a root base model to a model at the end of the ML supply chain (i.e., a leaf node with no dependents or outgoing connections).
Using the paths resulting from RQ1, we addressed RQ2 by analyzing the state of licensing in the ML supply chain, developing scripts and heuristics to look for parent/child license differences, and potential license compliance problems related to model/dataset dependencies. We define a parent/child license difference to refer to the difference between the license(s) of a parent and a child model/dataset. Assume, for example, that model $B$ is the child of model $A$. If model $B$ was licensed under MIT and model $A$ under Apache-2.0, we would say there is a parent/child license difference between $A$ and $B$. We avoid stronger terms such as “inconsistencies” and “incompatibilities”, since some of these differences may in fact be consistent and compliant with licensing terms. Two authors reviewed all distinct licenses in our dataset and grouped them into six categories, which can provide the basis for future research into license compatibility and licensing trends in the ML supply chain.
All scripts used for mining, data clean-up, and analysis can be found in our replication package (Stalnaker et al., 2025) and can be used to foster further research in this area.
4. RQ0: Documentation challenges encountered on Hugging Face
In this section, we outline and discuss documentation issues that were observed on the Hugging Face platform while mapping connections in the ML supply chain. Specifically, this preliminary RQ outlines the observed shortcomings and challenges to 1) elaborate potential pitfalls future researchers may face in mining Hugging Face and 2) to more thoroughly report documentation issues extant on Hugging Face that have been left unexplored by prior work.
Figure 2. Example of model gated behind Terms of Service requirement
4.1. Inability to Access Metadata
Not all model metadata was accessible through the Hugging Face API. There were 7,258 model cards (0.95%) that had to be scraped using a combination of Python’s requests (requests, 2024) and BeautifulSoup (beautifulsoup, 2025) libraries. The most frequent challenge (i.e., found in 99.4% of the cases) resulted from gating based on the acceptance of terms and conditions (see Figure 2). This was particularly the case for models owned by or relying on models created by larger entities such as Google or Meta. Twenty-two models and their metadata were gated behind age restrictions imposed by the model owner [^1] Throughout the paper, by ”model owner”, we refer to the entity that uploaded the model to Hugging Face. We make no claims as to whether that entity had the legal right to do so or if they are the original author/creator of the model., requiring a Hugging Face account to access. Accessing the full repository, and thus metadata, required logging in with a Hugging Face account and then, in most cases, providing an email address. This hurdle to even access the metadata for a model could potentially complicate the creation of fully automated analyzers and ML/AIBOM (Machine Learning/Artificial Intelligence Bill of Materials) generators for an ML supply chain.
4.2. Incomplete Metadata
We observed that the documentation provided by model owners on Hugging Face is often incomplete. This is a trend not just for smaller, infrequently used projects but also for large projects with tens of thousands of downloads. For example, OpenAI’s clip-vit-large-patch14-336 model, with 5,827,027 downloads, has a model card that is almost entirely incomplete, save for basic hyper-parameter and framework version information (OpenAI, [n. d.]). Only 37.8% of models and 27.6% of datasets declared any kind of licensing information in a machine-readable way. We discuss the prevalence of declared components further in Section 5.
Some models may provide information on training data and architecture by linking to scientific papers that include these details, but such a practice introduces an additional hurdle in easily and programmatically obtaining that data. Ultimately, the reason why many models do not even mention the datasets used would require further investigation, including interviewing/surveying developers, which is out of the scope of this work.
During the manual analysis of 100 model cards described in Section 3.3, we observed two examples where the Hugging Face team wrote documentation on behalf of model owners. The first instance was an image-to-text model uploaded by Microsoft with the disclaimer: “The team releasing TrOCR did not write a model card for this model so this model card has been written by the Hugging Face team” (Microsoft, nda). The second was a larger version of the same model, also owned by Microsoft (Microsoft, ndb). In both cases, the models were described by a research paper and released on GitHub. Understanding the criteria that the Hugging Face team uses to determine when and how to intervene with model cards requires further investigation, but we do note that according to a blog written by members of the Hugging Face team, they created/updated model cards in some instances to inform design decisions surrounding a new model card template (Ezi Ozoani, 2022).
4.3. Additional Metadata Issues
4.3.1. The ”unknown” license.
The Hugging Face platform’s documentation provides a list of recognized licenses each with a unique, standardized short-hand identifier that can be utilized by users when creating their model/data cards (Face, 2024a). This list includes common Open Source Software (OSS), Creative Commons (CC), and ML-specific licenses as well as an “Other” catch-all category which encompasses less common, custom, or modified licenses not included in Hugging Face’s predefined license list An “unknown” licensing option is also provided, but there is no guidance provided as to when this option should be selected. Future work will need to explore when and why model owners select this option, particularly given that acknowledging that the license is “unknown” suggests, at best, a lack of due diligence or understanding on the part of the owner or, at worst, that the model/dataset raises copyright infringement issues. We observed 4,419 (1.5%) models and 2,194 (4.5%) datasets that used this “unknown” license tag, which inherently raises compliance challenges for dependent models.
4.3.2. Naming problems.
Hugging Face users primarily use human-readable names, not unique identifiers, when supplying the metadata information for their base models/datasets. Unique identifiers do exist for all models and datasets on Hugging Face, but we did not observe any instances where this information was being used by model/dataset owners when referring to models or datasets they depended on. Instead, these references typically, but not always, followed the “owner/model” naming convention. For example, a model owner would refer to a base model as “FacebookAI/xlm-roberta-base” and not by its unique Hugging Face ID: 621ffdc036468d709f174364. This is problematic because these model references are not automatically updated when name changes occur. A change to the model, dataset, or owner name results in a new human-readable identifier that no longer matches the previous references to the model/dataset. As a result, documentation can easily become outdated, unhelpful, and potentially confusing.
We observed 596 models that had their human-readable identifiers changed. Instances such as 222gate/Blurdus-7b-v0.1 $\rightarrow$ gate369/Blurdus-7b-v0.1 involve a change to the name of the model owner. Other instances, such as aaditya/openbiollm-llama3-70b $\rightarrow$ aaditya/Llama3-OpenBioLLM-70B involve a change to the model name itself. In both cases, manual effort is required to track down and map the original references to the new names, which can make dependency management tasks increasingly difficult. Often the only way to determine the correct mapping is by relying on Hugging Face to redirect the page associated with a former name to the one associated with the most current name. This suggests that Hugging Face maintains some internal mapping, but it is unclear how long this mapping persists or what might happen in the event of a name collision.
Table 3. Top-10 Most Common Model Names in the Ecosystem
Model Name Count
ppo-LunarLander-v2 5,660
q-FrozenLake-v1-4x4-noSlippery 4,200
test 3,900
ppo-Huggy 2,684
dqn-SpaceInvadersNoFrameskip-v4 2,405
distilbert-base-uncased-finetuned-emotion 2,348
my_awesome_model 1,727
q-Taxi-v3 1,535
ppo-SnowballTarget 1,513
bert-finetuned-ner 1,297
Model names, without the additional owner information, are not guaranteed to be unique. Similar to how repository forks have the same name (but different owners) on GitHub, fine-tuned or forked models may also have the same name yet different owners on Hugging Face. For this reason, developers often specify models using the “owner/model” naming convention described previously. If this happens consistently, ambiguities can be avoided. However, we observed 16,477 cases where developers referred to base models only by the model name, excluding owner information. For example, a developer might use “roberta-base” to refer to “FacebookAI/roberta-base.” While it could be reasonable to assume that this shorthand refers to the popular model, this may not be the case: in our dataset, there were 39 other models also named “roberta-base,” all with different owners. Table 3 shows the top-10 most common model names in the Hugging Face ecosystem and their frequencies. If one of these models were referred to by name only, it would be nearly impossible to determine which model was being referenced. For example, more than 5,600 models share the model name “ppo-LunarLander-v2.” This further stresses the need for widespread adoption of unique model identifiers.
4.3.3. Missing or nonsensical references.
There were 34,159 instances where we were unable to map a declared dataset to a dataset found publicly on Hugging Face. These consisted of 4,755 unique declarations. It is impossible to know with certainty, but such declarations likely consist of datasets that have been removed or made private, are from external sources, are actually dataset descriptors, contain typos, or, in some cases, are Hugging Face usernames.
We were also unable to map 2,501 base model declarations to known IDs, representing a possible 1,371 unique missing models in total. As with the datasets, these models may have been removed, made private, renamed in a way difficult to trace, or may reference a model outside the ecosystem. It is also possible that developer typos result in the inability to identify the models.
In other cases, some of the references were nonsensical: in our analysis of the graph of all collected components in the ML supply chain (described further in Section 5), we detected 684 cycles in total, including 675 (98.7%) trivial cycles in which a model declares itself as a base model, despite the fact that a model should not be able to be its own ancestor. All of this further motivates the need to use unique and standardized identifiers when referring to models and datasets, as well as a need for validating reference information.
4.3.4. Models as datasets.
We observed 1,416 instances in which a model was listed as a dataset in the datasets field. In other words, a model reference was included in the metadata field specific to training datasets. Without more information, it is impossible to know whether this indicates a mistake on the part of the model owner or that the output of the declared model was somehow used for training.
4.3.5. Shortcomings of Hugging Face.
While base models are declared by some model owners, in many cases, the precise relationship between a derived model and its base model is left ambiguous, at least in the available metadata. Hugging Face does provide a separate field for architecture relationships, but there is no standardized way to specify situations involving fine-tuning, quantization, or using outputs for training. This ambiguity can have significant consequences since, particularly in the licensing context, the nature of the relationship is important in determining how licensing terms should be applied. For example, it is permissible to use a model licensed under the llama3 license for fine-tuning but impermissible to use its outputs to train a competing model (Meta, 2024). Additional fields specifying the nature of relationships between models would be useful, but their inclusion would also introduce additional overhead for model owners.
While uncommon and not a direct shortcoming of Hugging Face, we observed examples where model owners would declare the same dataset or base model multiple times (i.e., the same model/dataset had multiple entries in the respective metadata field). Specifically, 178 models declared a dataset at least twice, and 310 models declared a base model at least twice. One model, with otherwise robust documentation, declared the same base model 64 times. Again, while there are few instances of this duplication, they still serve as something of a canary in the coal mine, indicating a lack of validation on behalf of Hugging Face.
5. RQ1: The structure of the ML supply chain
Table 4. Top-10 Models by Likes
Model Likes License
runwayml/stable-diffusion-v1-5 10,987 creativeml-openrail-m
CompVis/stable-diffusion-v1-4 6,358 creativeml-openrail-m
stabilityai/stable-diffusion-xl-base-1.0 5,502 openrail++
meta-llama/Meta-Llama-3-8B 5,200 llama3
bigscience/bloom 4,647 bigscience-bloom-rail-1.0
meta-llama/Llama-2-7b 3,980 llama2
mistralai/Mixtral-8x7B-Instruct-v0.1 3,960 apache-2.0
stabilityai/stable-diffusion-2-1 3,778 openrail++
WarriorMama777/OrangeMixs 3,714 creativeml-openrail-m
meta-llama/Llama-2-7b-chat-hf 3,670 llama2
Table 5. Top-10 Models by Downloads
Model Downloads License
MIT/ast-finetuned-audioset-10-10-0.4593 677,161,532 bsd-3-clause
sentence-transformers/all-MiniLM-L12-v2 83,805,061 apache-2.0
facebook/fasttext-language-identification 53,517,573 cc-by-nc-4.0
google-bert/bert-base-uncased 52,724,375 apache-2.0
sentence-transformers/all-MiniLM-L6-v2 41,281,684 apache-2.0
openai/clip-vit-large-patch14 40,114,370 -
distilbert/distilbert-base-uncased 24,195,717 apache-2.0
openai/whisper-small 23,317,804 -
jonatasgrosman/wav2vec2-large-xlsr-53-english 18,212,315 apache-2.0
openai/clip-vit-base-patch16 16,999,162 -
Despite the challenges outlined in RQ0, we were still able to construct an ecosystem-level supply chain graph mapping the relationships between models and datasets. While we acknowledge that our supply chain graph represents only a subset of the total components in the full Hugging Face ecosystem, as noted previously, we are confident that we can still rely upon it to make observations about recurring trends and patterns.
To analyze the structure of the ML supply chain, we began by addressing the issues identified in RQ0 using different strategies depending on the issue. Where possible, we employed web-scraping techniques to pull metadata for models that were inaccessible through the Hugging Face API (Section 4.1). Models where even this analysis was impossible were excluded from the analysis. We standardized model/dataset names, preferring the unique Hugging Face ID (Section 4.3.2). Where naming ambiguities were present (multiple models with the same name, but different owners), we assumed the most popular model (Section 4.3.2). References that could not be mapped to existing models/datasets on Hugging Face were also excluded from analysis (Section 4.3.3). Models that were listed as datasets were excluded from the dataset analysis (Section 4.3.4).
5.1. Most popular models and datasets
Table 6. Top-10 Datasets by Likes
Dataset Likes License
fka/awesome-chatgpt-prompts 5,011 cc0-1.0
HuggingFaceFW/fineweb 1,539 odc-by
Open-Orca/OpenOrca 1,269 mit
OpenAssistant/oasst1 1,232 apache-2.0
gsdf/EasyNegative 1,128 other
Anthropic/hh-rlhf 1,092 mit
togethercomputer/RedPajama-Data-1T 1,025 -
Nerfgun3/bad_prompt 921 creativeml-openrail-m
tiiuae/falcon-refinedweb 765 odc-by
allenai/dolma 752 odc-by
Table 7. Top-10 Datasets by Downloads
Dataset Downloads License
hails/mmlu_no_train 16,146,384 mit
lighteval/mmlu 5,882,652 mit
argilla/databricks-dolly-15k-curated-en 4,671,897 -
ceval/ceval-exam 2,111,340 -
lavita/medical-qa-shared-task-v1-toy 1,792,638 -
lukaemon/bbh 911,736 -
haonan-li/cmmlu 633,608 cc-by-nc-4.0
chansung/requested-arxiv-ids-3 587,249 -
cais/mmlu 555,115 mit
allenai/ai2_arc 543,678 cc-by-sa-4.0
We begin with an overview of the landscape by describing the most popular models and datasets. We measure popularity in terms of engagement via the metrics provided by Hugging Face, likes and recent downloads. These metrics also serve our goal of capturing the current state of the ecosystem by highlighting what models and datasets are currently being engaged with. Tables 5 and 5 show the ten most engaged with models in our dataset by these metrics. Organizations such as OpenAI and Meta occupy several top positions. The models span a variety of use cases, including image and text generation.
Similarly, Tables 7 and 7 display the most engaged with datasets. The most common Task field is “Question Answering” for the most downloaded datasets and “Text Generation” for the most liked datasets. However, a variety of Data Modalities and Tasks are represented across both groups. For example, Data Modalities included both text and images, while Tasks included other specifics such as “Token Classification.” In fact, the most liked dataset by far, fka/awesome-chatgpt-prompt, is a repository of curated ChatGPT prompts rather than a training dataset for models, which suggests that it may be more prudent to examine download count to determine the most popular datasets. This is further supported by its correlation with the number of client projects on GitHub (Pepe et al., 2024).
5.2. Most depended upon datasets and base models
Our ability to evaluate the ML supply chain relies on the dependency information provided by model owners. On Hugging Face, this information is found in the model metadata, where model owners declare base models and datasets. Although many models on Hugging Face were not created independently, we note that only 117,245 out of 760,460 models (15.4%) declare any base model and, of these, most (111,568) declare only one base model. Consequently, the average number of declared base models is 0, as are the first quartile, median, and third quartile, as shown in Figure 3.
Figure 3. Models’ Number of Declared Base Models
Table 8. Top-10 Declared Base Models
Base Model Times Used Likes Downloads License
distilbert/distilbert-base-uncased 5,240 463 24,195,717 apache-2.0
stabilityai/stable-diffusion-xl-base-1.0 4,776 5,502 3,558,713 openrail++
runwayml/stable-diffusion-v1-5 2,438 10,987 3,790,115 creativeml-openrail-m
openai-community/gpt2 2,405 2,076 6,888,015 mit
unsloth/llama-3-8b-bnb-4bit 2,363 141 486,478 llama2
FacebookAI/xlm-roberta-base 2,169 526 6,196,387 mit
mistralai/Mistral-7B-v0.1 2,099 3,290 706,043 apache-2.0
google-bert/bert-base-uncased 1,576 1,665 52,724,375 apache-2.0
google-bert/bert-base-cased 1,472 235 5,191,167 apache-2.0
openai/whisper-small 1,469 180 23,317,804 apache-2.0
Table 8 lists the ten models that were most frequently declared as a base model by the models in our dataset, i.e., the models that were most frequently cited as dependencies by other models. They exhibit a varying number of likes and downloads and primarily come from large providers such as Meta, Google, and OpenAI. Notably, none of these models declare that they are derived from any base models. We observe some overlap with the ten most engaged with models by downloads, suggesting, not surprisingly, that popular models are also frequently used as base models.
In addition to base models, we investigate the datasets used for model training. We observe that 75,516 (9.9%) models declare at least one dataset. Of these, 64,343 (85.2%) declare a single dataset, 5,066 (6.7%) declare two datasets, and 1,637 (2.2%) declare three datasets. The number of datasets also exhibits long-tail behavior, with one model declaring 287 distinct datasets. However, the vast majority of models declare no dataset, with 0 declared models as the average and first three quartiles, as shown in Figure 4. We observe that this low percentage of declared datasets is particularly notable since, by their nature, nearly all ML models on Hugging Face should have been trained on at least one dataset, even if that dataset is not publicly available.
Figure 4. Models’ Number of Declared Datasets
Table 9. Top-10 Declared Datasets
Dataset Times Used Likes Downloads License
nyu-mll/glue 2,293 336 122,473 other
rajpurkar/squad 1,827 228 11,379 cc-by-sa-4.0
dair-ai/emotion 1,766 257 23,236 other
ilsvrc/imagenet-1k 1,470 345 17195 other
mozilla-foundation/common_voice_7_0 1,330 163 67,565 cc0-1.0
stanfordnlp/imdb 1,329 208 20,078 other
legacy-datasets/wikipedia 1,219 528 1,657 cc-by-sa-3.0, gfdl
google/xtreme 1,154 82 982
apache-2.0, cc-by-4.0, cc-by-2.0,
cc-by-sa-4.0, other, cc-by-nc-4.0
mozilla-foundation/common_voice_11_0 1,130 163 67,565 cc0-1.0
eriktks/conll2003 880 115 6,171 other
Table 9 describes the top-ten declared datasets. We observe no overlap with the top-ten most liked or downloaded datasets. This may be a consequence of the low declaration rate for datasets, a function of a dataset’s age, or the fact that popular “datasets” uploaded to Hugging Face may not be intended for model training, as observed above.
License declarations vary among the datasets. Half of the top-ten most declared datasets are made available under a license declared as “other.” One example from Google appears under six distinct licenses, including the Apache-2.0 license, several Creative Commons licenses, and an “other” license. While these are all listed in the dataset’s metadata, the dataset card’s field for licensing information simply says, “More Information Needed,” leaving it unclear as to how these licenses jointly apply to the dataset. This suggests that even very popular components can contain ambiguities that make the license compliance process more difficult.
5.3. Structure of the supply chain
5.3.1. Lengths of model supply chains
We define a lineage chain as a path from some root base model (i.e., one with no model dependencies) to a final sink node model (i.e., one without any model dependents). We examine 53,151 lineage chains for models that declare at least one base model, including cases where multiple chains lead to the same model. Of these, the average chain length to reach that model is 6.2 models. The most frequent chain length is of three models (12,480 chains), and, as shown in Figure 6, the first quartile is 3 models, the median is 4 models, and the third quartile is 9 models. The longest chain, beginning with cohereforai/c4ai-command-r-v01 and ending with Citaman/command-r-1-layer, contains 40 models. Excluding the first model, all models in this chain are owned by Citaman and incrementally count down from 39. That is, the models in the chain range from Citaman/command-r-39-layer down to Citaman/command-r-1-layer, suggesting that each model may represent an incremental improvement or an updated version.
Figure 5. Model Chain Lengths
Figure 6. Number of Licenses in Model Chains
5.3.2. Licenses present in model supply chains
There are typically relatively few parent/child license differences between models in these chains, with an average of 1.9 distinct licenses in a given chain, including unknown or undeclared licenses. As Figure 6 illustrates, the first quartile is just 1 license, and both the median and third quartile are 2 licenses. A chain most commonly contains 2 licenses (24,069 chains), and the highest number of licenses observed was 6, occurring in 17 different chains. A relatively small number of parent/child license differences simplifies the task of managing license compliance. However, these differences within a lineage necessitate examination of license compatibility, as well as a possible need to find mitigations for items in model chains when required by the license of a base model.
Figure 7. Number of Times the Owner of a Chain’s Final Model Appears in that Chain
5.3.3. Model ownership
We also observe that developers using Hugging Face often build off each other’s work rather than off their own models. The owners of nodes at the end of a chain appeared within that node’s chain only 1.4 times on average, showing that base models are frequently sourced from the community and supporting the idea that a common approach to training a model involves building off of previous work. In another interpretation, it could be that model trainers do not fill in base model information when building on their own work at least some of the time. Future work involving discussions with developers could shed light on the reasons behind this phenomenon.
In 82.9% of chains (43,507), the owner of the final model in the chain owned no other model in that chain. In the maximum case, the chain contained 39 models from the final model’s owner. Illustrating this further, as shown in Figure 7, the first quartile, median, and third quartile are all a single occurrence of the final model’s owner in a given chain: only the final model itself.
Based on our observations, a given owner publishes relatively few models to Hugging Face, owning, on average, 4 models in our dataset, though the data is skewed heavily toward relatively few highly-prolific owners. Of 190,136 distinct model owners in our dataset, 61.5% owned just one public model (117,016), while the most prolific account owned 4,610 models. The first quartile and median are both just 1 owned model, and the third quartile is 2 models. Notably, the large players mentioned above, including Meta and OpenAI, are not present in the top-10 most prolific model owners: those spots instead go to smaller stakeholders in the AI market. The top-ten model owners on Hugging Face are shown in Table 11 and the top-ten dataset owners in Table 11. There is only one overlap between the two lists.
Table 10. Top-10 model owners on Hugging Face
Owner Model Count
mradermacher 4610
thebloke 3863
huggingtweets 3830
tkcho 3810
jeevesh8 3723
prunaai 3411
lonestriker 3241
sail-rvc 3182
larryaidraw 3046
cyberharem 2978
Table 11. Top-10 dataset owners on Hugging Face
Owner Dataset Count
open-llm-leaderboard-old 7059
cyberharem 3384
autoevaluate 3232
recag 2345
liuyanchen1015 1558
adapterocean 810
tyzhu 807
joey234 705
lots-of-loras 500
results-sd-v1-5-sd-v2-1-if-v1-0-karlo 462
6. RQ2: licensing of models and datasets on Hugging Face
The majority of models (62.2%) and datasets (72.4%) did not declare licensing information in a machine-readable way (see Section 3.3). To better illustrate how models and datasets are licensed, we organize the licenses that we observed into six categories. The categories are intended to be indicative of the origin or purpose of the licenses (whether the licenses were designed for open source software, for ML components specifically, or other purposes) and are not intended to convey any legal characteristics, particularly given the uncertainty involved in applying these licenses in the ML context. We define the following categories:
  • OSS: open source software licenses such as MIT, Apache, and GPL
  • CC: Creative Commons licenses
  • ML: ML-specific licenses such as open-rail and llama
  • Data: licenses that are intended to specifically protect data such as odbl and c-uda
  • Other: the “other” license category on Hugging Face
  • Unknown: the “unknown” license category on Hugging Face
6.1. Most common licenses
Table 12. Top-10 licenses for models.
Models
License Type Count
Percent
(of licensed)
apache-2.0 OSS 119,449 41.6%
mit OSS 51,184 17.8%
openrail ML 30,095 10.5%
creativeml-openrail-m ML 20,396 7.1%
other Other 16,447 5.7%
cc-by-nc-4.0 CC 8,318 2.9%
llama2 ML 6,158 2.1%
unknown Unknown 4,419 1.5%
cc-by-4.0 CC 4,269 1.5%
llama3 ML 3,335 1.2%
Table 13. Top-10 licenses for datasets.
Datasets
License Type Count
Percent
(of licensed)
mit OSS 13,501 27.9%
apache-2.0 OSS 12,372 25.6%
openrail ML 5,501 11.4%
cc-by-4.0 CC 2,673 5.5%
unknown Unknown 2,194 4.5%
other Other 1,837 3.8%
cc-by-sa-4.0 CC 1,256 2.6%
cc-by-nc-4.0 CC 1,014 2.1%
cc-by-nc-sa-4.0 CC 946 2.0%
cc0-1.0 CC 799 1.7%
There were 71 unique declared licenses for models and 70 for datasets. Tables 13 and 13 provide an overview of the most frequently observed licenses.
We classify licenses by the classes defined above. A breakdown of the prevalence of these classes can be found in Table 14. Both models (62.5%) and datasets (58%) are most likely to be licensed under approved OSS licenses. This is particularly interesting for datasets, where we would suspect to see mostly CC or Data licenses, as datasets are not “software” in a strict sense of the word. This may reflect that Hugging Face developers view models and datasets more similarly to programs than to data or that developers rely on OSS licenses because they are already familiar with their terms. However, the relationships between these preexisting licenses and these new types of components in the ML supply chain, as well as the implications of licensing such components with such licenses, are currently unknown. Further research is needed to understand the reasoning behind these licensing choices, as well as to understand the interactions between many of these licenses and ML components.
Table 14. Prevalence of License Classes
Models Datasets
License Class Count Percent License Class Count Percent
OSS 179,421 62.5% OSS 28,047 58%
ML 66,540 23.2% CC 9,151 18.9%
CC 19,856 6.9% ML 6,481 13.4%
Other 16,447 5.7% Unknown 2,194 4.5%
Unknown 4,419 1.5% Other 1,837 3.8%
Data 326 0.1% Data 559 1.2%
Combination 134 0.05% Combination 87 0.2%
Total Licensed 287,143 100.0% Total Licensed 48356 100.0%
Both models and datasets are distributed with (i) OSS licenses specific to software (both restrictive and permissive), (ii) Creative Commons (CC) licenses, that are not software-specific yet are being used for software (e.g., Stack Overflow’s adoption of CC licenses), and ML-specific licenses. For models, our findings are consistent with those of Pepe et al. (Pepe et al., 2024), showing that permissive OSS licenses such as Apache-2.0 and MIT are popular in this space. This is not surprising, as such licenses are also among the most popular for open-source projects (Vidal, 2023; Vendome, 2015) and allow commercial/closed-source exploitation. We note that, in general, the top licenses adopted on Hugging Face are traditionally understood to be more permissive, even if their exact application in this novel context is not fully understood.
We also note the adoption (for models, but especially for datasets) of different variants of the CC license, including a permissive one (CC-BY), restrictive ones (CC-BY-SA and CC-BY-NC-SA), and licenses limiting non-commercial use only (CC-BY-NC).
Besides OSS licenses, we see a large proportion of different kinds of ML-specific licenses. These include both licenses originating from open-source initiatives (openrail, creativeml-openrail-m), and licenses originating from companies (e.g., Meta’s llama2 and llama3 licenses). Such licenses tend to differ from typical redistribution terms established by software licenses by, for example, introducing “behavioral” constraints; requiring responsible model usage; and prohibiting adoption for harmful or unethical uses or uses that do not take models’ limitations into account. Likewise, the llama2 and llama3 licenses impose limitations related to commercial use, use in products or services having more than 700 million monthly active users, and use to train other models, unless such models are redistributed, under the same licenses, as derivative works of llama models.
Last but not least, we note the increased prevalence of Hugging Face’s “Other” license which was ranked 7th in Pepe et al. (Pepe et al., 2024) but has moved to the 5th spot since. In fact, 3.8% of datasets and 5.7% of models are licensed with some “Other” license. Prior work has found that such license proliferation can make license compliance tasks more time-consuming and difficult, since compliance teams are no longer dealing with known quantities (Wintersgill et al., 2024).
6.2. Parent/child license differences
Table 15. Parent/child license difference categories and their frequencies
Difference Type Description Example Count Percentage
No Overlap
Both the parent and child are licensed, but
there is no overlap in their licensing
cc-by-nc-4.0 $\rightarrow$ apache-2.0 36,441 54.8%
Licensing Dropped
The parent is under one or more licenses,
but the child has no specified license
apache-2.0 $\rightarrow$ UNDECLARED 19,435 29.2%
Licensing Added
The child is under one or more licenses, but
the parent has no specified license
UNDECLARED $\rightarrow$ apache-2.0 10,548 15.9%
License(s) Added
In a multi-licensing scenario, the child retains
all of its parent’s licenses and adds one or more
additional licenses.
apache-2.0 $\rightarrow$ apache-2.0 |||| mit 28 0.04%
License(s) Removed
In a multi-licensing scenario, the child is no
longer under one or more licenses that apply
to its parent, but is still licensed and adds no
additional licenses.
apache-2.0 |||| bsd-3-clause $\rightarrow$ bsd-3-clause 8 0.01%
Total Changes - - 66,460 100.0%
Table 16. Most common parent/child license differences
Parent License Child License Count Percentage
cc-by-nc-4.0 apache-2.0 11,001 30.19%
apache-2.0 cc-by-nc-4.0 7,427 20.38%
cc-by-4.0 cc-by-nc-4.0 2,626 7.21%
other llama3 1,720 4.72%
cc-by-nc-4.0 cc-by-4.0 1,444 3.96%
other apache-2.0 1348 3.70%
apache-2.0 mit 1,241 3.41%
cc-by-nc-nd-4.0 cc-by-nc-4.0 1,180 3.24%
llama3 other 1,083 2.97%
apache-2.0 cc-by-4.0 782 2.15%
Our dataset contains 274,104 distinct parent/child relationships. We observed 66,460 instances (24%) where the licensing of a child model (i.e., derivative model) was different from that of its parent (i.e., base model). Table 15 provides an overview of the types of differences observed. In nearly a third of cases, child models specified no licensing information despite information being available for their parent(s), leading to a potential compliance violation. We also observe that once license information has been dropped, it is unlikely to be restored by later links in the chain, with license restoration behavior observed in only 15.9% of instances. A complete shift in licensing between parent and child was seen in 54.8% of cases (i.e., no licenses were shared between the two). Of these shifts, 18.5% involved the “unknown” or “other” license category. Table 16 shows the top ten such shifts, and Figure 8 provides more information on how licensing decisions shifted across class boundaries between parents and their children. While ML-specific licenses seem a good fit for models, model owners may be opting for licenses they are more familiar, and comfortable with (OSS for software developers and CC for data scientists). We observe that most shifts occur across the OSS and CC boundary (68.3%), perhaps suggesting some tension between the licensing preferences of those with data science and software development backgrounds. Additionally, depending on the specifics of the situation and the licenses involved, these shifts can also potentially introduce license incompatibilities, such as dropping non-commercial requirements or failing to apply the requirements of a copyleft license.
Figure 8. Parent/child license differences across class boundaries (Total differences: 29,708)
6.3. Dataset license and model license combinations
Table 17. Model license and dataset license pairs
Model License Dataset License Count Percentage
apache-2.0 other 8196 18.9%
apache-2.0 cc0-1.0 4621 10.6%
mit mit 4479 10.3%
apache-2.0 apache-2.0 3916 9.0%
apache-2.0 unknown 2993 6.9%
apache-2.0 mit 2734 6.3%
mit other 2290 5.3%
apache-2.0 cc-by-4.0 2076 4.8%
apache-2.0 cc-by-sa-4.0 1934 4.5%
mit apache-2.0 1739 4.0%
The licenses of datasets that are used during training can have an impact on the licensing decisions for the final ML model. Here we consider the pairings of dataset licenses and the resultant model licenses. There are 623 distinct model/dataset license combinations within our dataset across 43,455 such pairings. (The top ten most frequent can be found in Table 17.) However, since we (and Hugging Face) aggregated all “other” licenses and treated them as one quantity, these numbers likely overestimate the consistency in the space. In 41 of the 623 combinations, the license of the model exactly matches the license of at least one dataset it was trained on. The most common combination is a model licensed under the apache-2.0 license and a dataset under a custom or “other” license. In total, 11,731 (27%) pairs involve a dataset under a custom or “other” license. This can make using these datasets and models problematic since retrieving and evaluating the licensing terms can be difficult and time-consuming.
6.4. Other licensing findings
6.4.1. Multi-licensing.
A few models (188) and datasets (104) were released under more than one license. Multi-licensing is an existing phenomenon in the world of open-source software but is made more complex in the ML model context by the presence of novel license combinations that are not yet well understood. Examples include the 48 models we observed multi-licensed under apache-2.0 and cc-by-nc-4.0 as well as models under both OSS and ML licenses, such as MIT and OpenRAIL. We observe examples of such multilicensing in the most declared datasets in Table 9. We also observed datasets that were released under as many as six distinct licenses. The Hugging Face metadata provides no way of determining the relationship among these licenses, which could be either an AND or an OR. This distinction is critical to determining compliance. For example, if a user who uploaded a llama2 derivative model chose to make it available with an OR relationship between the apache-2.0 and llama2 licenses, this would contradict the exclusive llama2 licensing of its base model, resulting in a violation that could propagate to models further down the chain. While we observed only a few instances of multi-licensing, such activity motivates the need to understand the interplay between the various license classes and for Hugging Face to supply a standardized way to specify the intended relationship between licenses.
6.4.2. Naming requirements.
Some licenses, like Llama3, impose a generational limitation (Laurent, 2004) that specifies naming requirements for derivative works. According to the license, models that are built on a model under this license must have names prefixed with “llama3” (Meta, 2024). In our dataset, there are 3,335 models licensed under llama3. However, of these, only 473 (14.2%) models correctly have names beginning with “llama3” as required by the license. Only 768 (23%) models even contain the term “llama3” in their name. Just 2,421 (72.6%) models contain the term “llama” at all, and 384 (11.5%) only contain the string “l3.” The problem is likely even worse in practice, as these statistics only include models that documented that they were under the llama3 license.
6.4.3. Attribution.
As was observed during the manual model card analysis described in Section 3.3, attribution appears to be important to model owners regardless of licensing status: rather than requesting or requiring attribution through a license, even models without a license often asked for some form of attribution (29/84), typically in the form of citations. In fact, 83,628 models (11.1%) mined using the Hugging Face API included a link to a paper on arxiv (arxiv, [n. d.]) in their tags. In our manual analysis of model cards, we observed that the model owner was typically also affiliated with (e.g., an author on) the linked research paper. This desire for recognition is reflected by the top licenses in the space, namely MIT and Apache-2.0, each of which specifies attribution requirements. (A similar historical example is the evolution of Creative Commons licenses in the early 2000s, where the overwhelming preference for attribution led to its becoming a default term (glenn, 2004).) Given the research community’s adoption of Hugging Face, it may be worthwhile to add dedicated metadata fields to facilitate the standardization of citation information.
7. Discussion
In this section we discuss the implications of our findings and potential directions for future work.
7.1. License compliance in the ML supply chain
In Section 5, we showed that the structure of the ML supply chain consists primarily of relatively short component chains. This presents a radically different landscape from the traditional software supply chain, in which a single piece of software can depend on potentially hundreds or even thousands of components. However, just because fewer dependencies are declared does not mean that ensuring license compliance is any easier in this ecosystem. In fact, as we observed, licensing information is often missing (Section 6) and there can be licensing differences between parent and child models in a model’s lineage (Section 6.2). Compounded with our observation that datasets may also have unclear licensing (Section 6.1) and may go undeclared by models (Section 5.2), the state of documentation on Hugging Face appears to be counterproductive to license compliance tasks.
The task of managing different licenses across a model’s components presents a novel challenge due to the advent of ML-focused licenses and a heavy reliance on CC licenses. While previous work has investigated compatibility between various traditional software licenses (Kapitsaki et al., 2022; Liu et al., 2024), new licenses, such as the OpenRAIL family and the Llama licenses, require analysis to understand how they interact with existing licenses and with each other. Additionally, ML-focused licenses also indicate legitimate uses for the model and any derivatives, further separating them from the traditional open source definition. Proponents of OpenRAIL argue that existing open-source licenses are “ill-adapted” to handle ML components’ differences from source code, particularly with respect to the responsible use of such components (Ferrandis, [n. d.]). This necessitates future work to provide a detailed analysis of software licenses with respect to ML components, including an analysis of these licenses and the ways in which licenses interact with ML components differently from traditional software.
It also becomes crucial to educate model owners and users on the terms associated with licenses. For example, we observed that the simple naming requirements imposed by the Llama3 license were followed in only 14.2% of cases (Section 6.4.2). This suggests either extensive noncompliance or, more likely, an ignorance of the terms being violated. To concretely identify the cause and provide potential mitigations, future work should investigate developers’ views and understanding of licenses in the ML supply chain.
It also remains unclear, particularly with the relationship between datasets and models, how exactly license terms, such as GPL’s copyleft, should be applied. To what extent should a model be considered a derivative work of the dataset(s) on which it was trained? More concretely, if a model trainer trains a model on a dataset governed by a copyleft license and then distributes the model, but not the dataset, has the trainer violated the dataset license? These questions and many others will likely need to be resolved by legal teams, regulators, and potentially the courts.
Lastly, in Section 6.4.1, we also observed instances of ambiguity introduced through multi-licensing. Determining the intended relationship among the multiple licenses assigned by the model/dataset owner can be a difficult task, and different pitfalls come with the assumption of an OR versus an AND relationship. Model hubs should include standardized ways for model/dataset owners to express these relationships and educate users on why they might want to use a multi-licensing scheme. It is left to future work to explore why some developers are presently releasing their models/datasets under multiple licenses.
7.2. Potential difficulties in using model cards as ML/AIBOM
Amidst the growing push for better transparency and security in software through Software Bills of Materials (SBOMs) (Xia et al., 2023a; Stalnaker et al., 2024), calls have been made for similar BOMs for ML components (Stalnaker et al., 2024; Xia et al., 2023a; Barclay et al., 2019, 2022). These documents serve as inventories of all components within a piece of software and can document a variety of the software’s traits, including dependencies, licensing, and security. Distinct ML/AI Bills of Materials (ML/AIBOMs) may be necessary to address such components’ different inputs, security concerns (such as model poisoning), and ethical considerations (Xia et al., 2023a, b; Bi et al., 2024; Lu et al., 2024). The SPDX (spdx, 2025) Working Group, operating under the auspices of the Linux Foundation, is developing guidelines and a proposed standard for such an ML/AIBOMs (Bennet et al., 2024) that will complement the SPDX ISO Standard ISO/IEC 5962:2021, which describes the use of SBOMs to document the components used in creating a software system (ISO/IEC 5962:2021, 2021). CycloneDX (cyclonedx, 2025a) is also working on support for ML/AIBOMs (cyclonedx, 2025b).
Prior work has also suggested that model and data cards can serve as an ML/AIBOM (Stalnaker et al., 2024), or the information they provide can assist with their creation (Xia et al., 2023b; Pepe et al., 2024), but our work suggests that these tools are not yet robust enough to serve this purpose. As we highlight in Section 4, the information provided by model cards is often missing key elements, including datasets that the models were trained on. This immediately obviates many of the benefits of ML/AIBOMs, including understanding licensing obligations that might be associated with that dataset, being aware of potential model poisoning attacks, and providing the ability to select models trained on ethically sourced data. Additionally, the cases where model cards were locked behind Terms of Service agreements or other restrictions (Section 4.1) could further limit their usefulness as ML/AIBOMs to consumers. Our work provides additional evidence that, in practice, model cards contain little actionable information, making them difficult to use as ML/AIBOMs (Bhat et al., 2023).
7.3. The need for follow up surveys and interviews
Our study raises many questions concerning what motivates developers when they make decisions concerning documentation, model reuse, and licensing. Conversations with Hugging Face developers through surveys and interviews can help to bridge this gap. There are numerous key questions such work should consider. What factors influence developers when choosing a license for their dataset/model (Section 6.1)? Why do developers so commonly opt to use an “other” license (Section 6.1)? What motivates developers to multi-license a model/dataset and what is the intention behind the license interaction (Section 6.4.1)? Why do so few developers fine-tuning the llama3 model comply with the naming requirements imposed by the llama3 license (Section 6.4.2)? What is the primary motivation for developers to write documentation for their models? For each of the various stakeholders in the space (model owners, fine-tuners, and users), which metadata fields do they perceive as most important? Why are some models listed as datasets (Section 4.3.4)? Answers to these questions will inform how future tools, practices, techniques, and even educational materials should be developed to have maximal impact on the emerging ML community.
7.4. The addition of fields for recording model relationship information
As the relationships between ML models become more complex, it becomes increasingly important to keep a record of those relationships (e.g., fine-tuned, quantized, ensemble, student/teacher, etc.). This information is crucial not only for transparency and understanding how the model was created, but also for license compliance. Just as in traditional software, licenses may have terms that govern the manner in which the component may be (re)used and distributed. For example, many generative models disallow using generated content to train a derivative model (openai-terms, 2023). As noted in Section 4.3.5, Hugging Face provides no standardized way to express relationship information. This inability to specify relationships may have resulted in some model owners declaring models used in student-teacher relationships as datasets (Section 4.3.4). We suggest that model hubs provide a standardized, machine-readable way of conveying this type of information. Just as Hugging Face currently provides a standard list of architectures, it could also provide a curated list of relationship types.
7.5. The need to use unique identifiers
While unique identifiers already exist on Hugging Face, they are not currently being used by model/dataset owners. This can result in problems, particularly when trying to trace model/dataset provenance. Model hubs should encourage their users to use these existing unique identifiers, perhaps by including tools that facilitate the creation of model cards or the provision of metadata. For example, a developer could supply a colloquial name for a model hosted on the platform, such as the familiar owner/model convention, and the tool could automatically conduct a search and, where possible, map the provided name to the unique ID. This would immediately alert the developer of typos, remove ambiguities, and also prevent the name drift issues we discussed in Sections 4.3.2 and 4.3.3. Furthermore, such a system could potentially allow for automatic updates (or request manual developer updates) when the information for downstream elements changes, thereby facilitating an ecosystem which is more supportive of supply chain management tasks and lifting the burden from downstream developers.
7.6. Differences between the ML supply chain and traditional software supply chain
Our findings highlight a notable difference between the ML supply chain and the traditional software supply chain. In many software ecosystems with dependency management, dependency information for components in the software supply chain is organized in manifest files such as pom.xml, package.json, or requirements.txt. These files can then be used to generate other dependency tracking documents like SBOMs and have also been used by forges like GitHub to build dependency graphs (dep, [n. d.]). Notably, however, these manifest files have utility beyond dependency tracking. They are necessary for setting up a fresh installation of the software – that is, the list of dependencies must be downloaded in order for the software to function properly. This is not the case with the information supplied by developers in model cards. Because erroneous mappings to datasets and base models do not leave the model in question unusable as a practical matter, it is therefore easier for typos and other mistakes to go unnoticed, and thus uncorrected. This disconnect between correctness and functionality necessitates tooling that has similar functionality and richness to dependency management tools for traditional software, as it leaves ample opportunity for errors without comparable means of checking for and correcting them.
7.7. Quality control for model hubs
Our analysis of the model documentation available on Hugging Face shows that the information that developers provide is often incomplete, malformed, or only available in human-readable text. We observed some documentation issues, such as the presence of duplicates in lists of datasets and base models (Section 4.3.5) or the declaration of a model as its own base (Section 4.3.3), that could be easily corrected by automated tools like linters. Additionally, an automated tool could also provide standardization for data fields to facilitate metadata analysis. For example, the variants for how one can specify that there are no datasets or base models ([], “”, and None) should be standardized to a single type. We are not suggesting that model hubs should be responsible for curating the metadata of all models on their platforms, but where automated checks can efficiently be put into place, they should be employed.
Potential solutions could closely resemble approaches to automatically check the quality of bug reports. Similarly to how tools can check for important components of a bug report (e.g.,, expected behavior, steps to reproduce, etc.(Chaparro et al., 2017), in theory, tools may be able to check for missing information in model cards. Some work is already emerging in this area, such as DocML (Bhat et al., 2023) and AIMML (Tsay et al., 2020), however, more work needs to be done to understand what suggestions should be made to model owners, how to handle complex licensing scenarios, and how to best implement tools that assess ML documentation.
7.8. Potential pitfalls when mining Hugging Face
The numerous documentation challenges that we discuss in Section 4, demonstrate that researchers relying on Hugging Face mining to collect information on the ML supply chain must proceed thoughtfully, with caution and care to avoid potential pitfalls. Unlike source code forges like GitHub, it seems that some model owners may be uploading and publishing models as a by-product of their workflows. As such, we speculate that some models are made publicly available not because the owner intends them to be reused by the community, but rather because the owner is using the resources of Hugging Face, Google Colab, or some other cloud provider. (This might explain why we observed so many models with the same names. See Section 4.3.2.) More research is required to understand how such tools might populate the supply chain with such “intermediate” models, or those that are otherwise not intended for public consumption. To mitigate the risk such models pose, future studies should be mindful of the models they include in their analysis, erring toward those that are more popular. However, determining popularity may not be as easy as sorting by downloads or likes. As shown in Table 7, the most liked “dataset” on Hugging Face is in reality a repository curating ChatGPT prompts. Future studies will be left to explore the prevalence of curated repositories hosted as datasets and the exact nature of potentially ad hoc developer workflows.
8. Threats to Validity
Construct Validity. We rely on information that is available through the Hugging Face API in a machine readable format. We addressed cases where information was missing from this source by applying Python’s request and BeautifulSoup libraries to retrieve missing information. However, this means that we likely missed licensing and dependency information that was presented in natural language. We measure and mitigate this risk with a review of the top-100 “unlicensed” models sorted by number of downloads, which showed that the majority did not, in fact, declare a license or declared a license in a non-standard or transitive way.
Internal Validity. We followed best practices in our mining and data cleaning approaches. Conclusions were drawn based on empirical evidence found in the data. We relied on tags applied to components to detect their licenses, but such information might be present in other, human-readable forms or through links to other resources, such as a research paper or another repository. The labeling of license categories is based on our understanding of the licenses. To mitigate bias in these labels, two researchers examined each license to confirm the label applied to it.
External Validity. The generalizability of our findings is restricted by our choice of Hugging Face, which might not be fully generalizable to all model-sharing platforms. It is possible that we missed other challenges and obstacles faced by other model hubs or that different model-sharing sites have different levels of documentation and dependency management techniques. Additionally, given that the field of AI is rapidly evolving and new models are created every day, the conclusions we glean may be different if we were to take an updated snapshot of the supply chain at some future point. The extent to which this factor might change our results is unknown, but we maintain that it is important to capture an understanding of this nascent supply chain so that future investigations can understand the ways in which the supply chain has changed over time.
9. Conclusion
In this work we conduct a comprehensive mining study and analysis of the ML supply chain on Hugging Face. We construct a reusable dataset of over 750,000 models and 175,000 datasets, which we make freely available in our replication package (Stalnaker et al., 2025). Using this data we create a supply chain graph that models the relationships between ML models and the models they depend on, describe the notable features and properties of this supply chain, analyze the current state of documentation on Hugging Face, and consider the current state of licensing and license compliance in the ecosystem. We offer a rich discussion of the challenges that not just Hugging Face but any sizable model hub will encounter. Lastly, we offer concrete and actionable recommendations that can be adopted by any model hosting platform to facilitate better supply chain management.
Data availability
We provide an anonymized replication package containing the raw mined data, our cleaned and normalized dataset, scripts for data cleaning, mining utilities, analysis scripts, and other data required for verifiability (Stalnaker et al., 2025).
10. Acknowledgments
This research has been supported in part by NSF grant CCF-2217733. Any opinions, findings, and conclusions expressed herein are the authors’ and do not necessarily reflect those of the sponsors. A complete, detailed list of image attributions can be found at (Stalnaker et al., 2025).